# Yi-Fang Wang, Chia-Ming Lee, Yi-Shiuan Chou

### v2

@@ Abstract
As DeepFake video manipulation techniques escalate, posing profound threats, the urgent need to develop efficient detection strategies is underscored. However, one particular issue lies with facial images being mis-detected, often originating from degraded videos or adversarial attacks, leading to unexpected temporal artifacts that can undermine the efficacy of DeepFake video detection techniques. This paper introduces a novel method for robust DeepFake video detection, harnessing the power of the proposed Graph-Regularized Attentive Convolutional Entanglement (GRACE) based on the graph convolutional network with graph Laplacian to address the aforementioned challenges. First, conventional Convolution Neural Networks are deployed to perform spatiotemporal features for the entire video. Then, the spatial and temporal features are mutually entangled by constructing a graph with sparse constraint, enforcing essential features of valid face images in the noisy face sequences remaining, thus augmenting stability and performance for DeepFake video detection. Furthermore, the Graph Laplacian prior is proposed in the graph convolutional network to remove the noise pattern in the feature space to further improve the performance. Comprehensive experiments are conducted to illustrate that our proposed method delivers state-of-the-art performance in DeepFake video detection under noisy face sequences. The source code is available at https://github.com/ming053l/GRACE.
@@ 1 Introduction
DeepFake image and video manipulation techniques have emerged as the most well-known forgery generation applications, with far-reaching impacts on numerous individuals. Generally, facial manipulation schemes can be classified into four categories [1]: 1) entire face synthesis, 2) attribute manipulation, 3) identity swap, and 4) expression swap. Identity swap schemes have the most significant impact as they can be used to fabricate fake news targeting specific politicians. Many DeepFake detection techniques focus on identifying such fake videos using supervised learning methods with a pre-collected large-scale training set [5, 6, 7, 8, 9, 10, 11, 12].
Several advanced learning strategies have been proposed to enhance the performance of DeepFake image detection. For instance, methods in [7, 8] treat DeepFake image detection as binary classification tasks. The method in [10] introduces a novel multi-task learning approach to improve robustness and effectiveness. The authors in [9] also assert that traditional convolutional neural networks (CNNs) can be used to easily extract fake traces. However, the generalizability of such supervised learning strategies may be limited, as it is challenging to recognize DeepFake images generated by unknown GANs [13] due to difficulties in discerning out-of-distribution feature representations. To address the generalizability issue, semi-supervised learning is considered in [13, 14, 15] to capture common fake features from selected representative GANs, assuming that most GANs might share similar identifiable clues. Pairwise learning is then employed to learn these common features from the training set, improving generalizability for DeepFake image detection [14, 13]. Additionally, knowledge distillation is proposed for DeepFake detection by effectively transferring the weights of complex models to smaller models for enhanced generalizability [11].
In the domain of DeepFake video detection, numerous sophisticated approaches have recently emerged [16][17][18][19][20][21] [22][23][24][25]. On the one hand, these methods extend DeepFake image detection techniques by averaging the predictions of individual frames to assess a video’s authenticity [18][23][24][25]. On the other hand, the temporal inconsistency feature is exploited for DeepFake video classification using supervised learning approaches, as demonstrated in [20][21][16][17]. Specifically, state-of-the-art DeepFake video detection primarily focuses on exploiting various priors, such as bio-informatics clues [24], facial warping artifacts [25], and noise patterns [18]. Recently, several advanced techniques have been proposed to enhance DeepFake video detection performance. CORE [26] introduces a novel approach for learning consistent representations across different frames, while RECCE [27] employs a reconstruction-classification learning scheme to capture more discriminative features. DFIL [28] proposes an incremental learning framework that exploits domain-invariant forgery clues to improve generalization ability. TALL-Swin [29] utilizes a thumbnail layout and Swin Transformer to learn robust spatiotemporal features for DeepFake detection. UCF [30] focuses on uncovering common features shared by different manipulation techniques to enhance generalizability.
Another strategy to effectively detect the DeepFake image/video could be the extra-clue-inspired approach. In [24], a novel bio-feature—the Photoplethysmography (PPG) response—is utilized to differentiate DeepFake videos, as real and fake videos exhibit distinct PPG features. A critical limitation is the necessity for high-resolution videos and images to effectively capture PPG cues. Moreover, [25] investigates warping artifacts at the boundaries of DeepFake videos, which arise due to the limited resolution of synthesized facial components. However, contemporary GANs can generate high-resolution, realistic faces, rendering the resolution-inconsistency clues in [25] potentially less significant. Similarly, Face X-ray [18] leverages the boundary between real and fake facial regions as features, positing that the noise patterns of these parts differ and enabling traditional deep neural networks to identify DeepFake videos. As Face X-ray [18] demonstrates superior performance and robustness, recent DeepFake video detection techniques concentrate on uncovering more reliable signatures produced by GANs to enhance detection performance. Concurrently, [16] introduces deep Laplacian of Gaussian and the loss of isolated manipulated faces to bolster the generalizability of DeepFake video tasks.
Recent research has concentrated on developing robust DeepFake detection models for compressed videos, as seen in [20][31][32]. The frequency component analysis method is employed to uncover intrinsic features and enhance performance under compression settings [31]. However, frequency-aware features may prove ineffective under high compression (with high-frequency reduction) or noisy conditions (with high-frequency amplification). The $F^{3}$-net [32] selects two complementary frequency bands as clues, devising a novel network to learn frequency-aware features that reveal subtle forgery artifacts. Specifically, Frequency-aware Image Decomposition (FAD) is designed to learn subtle forgery patterns, while Local Frequency Statistics (LFS) primarily extracts high-level semantic features. This approach improves performance for low-quality inputs. A recent development in DeepFake video detection, the Anti-DeepFake Transformer (ADT), is proposed in [19], with robustness confirmed through cross-dataset evaluation. Recent studies [33][34][35] have highlighted the vulnerability of DeepFake detectors to adversarial perturbations. Therefore, adversarial defense with DeepFake detection, such as [36], have been attracting recently. In [36], it shows the effective solution by leveraging the statistical inference on the CNNs for achieving better robustness to adversarial examples.
@@
Figure 1: Example of the detected faces from two videos using RetinaFace (top) and Dlib (bottom).
@@
Even video compression could reduce the detection rate of the face detectors. We randomly select 200 videos from the FF++ dataset [1], featuring varying compression ratios (raw, c23, c40), and extract 16 frames from each video for face detection analysis. We employ state-of-the-art face detection tools, such as RetinaFace [42] and Dlib [43], to substantiate our observations. Table I presents the face detection outcomes for the sampled videos. Notably, the predicted probability of 8 and 2 videos using RetinaFace [42] and Dlib [43] falls within the $[0,0.33]$ range, respectively, implying there are 8 and 2 facial images are mis-detected. Additionally, for uncompressed videos, 19 videos exhibit accuracy lower than $66\%$, highlighting the imperfections of face detectors. The question is raised: Could the current DeepFake detection methods be robust to such noisy face sequences? The answer is negative. We simply replace the $40\%$ facial images with background ones for the testing set of FF++ [1] with raw setting and evaluate the performance using Xception [6]. Unsurprisingly, the accuracy dropped significantly after replacement. An effective solution to deal with the issues raised by noisy face sequences for DeepFake video detection is highly desired.
Figure 2: Flowchart of the proposed GRACE with Graph Laplacian regularizer for robust DeepFake video detection.
@@
In light of the escalating threat posed by various malicious attacks on face detectors that aim to undermine their reliability, this paper presents a pioneering Graph-Regularized Attentive Convolutional Entanglement (GRACE) with Laplacian Smoothing learning approach. GRACE leverages contextual features in both temporal and spatial domains to effectively detect DeepFake videos under noisy face sequences. We meticulously incorporate sparsity regularization into our model to prioritize the features of valid face images within the noisy face sequence. By employing the proposed Feature Entanglement (FE) technique, an affinity matrix is constructed to amalgamate the spatiotemporal features, ensuring that each node possesses at least one feature descriptor originating from valid face images. Ultimately, Graph Laplacian (GL) smoothing regularization is ingeniously integrated into the Graph Convolutional Network (GCN) to further suppress noisy nodes, thereby significantly enhancing the performance of DeepFake video detection. The main contributions of this paper are three-fold:
@@
  • We propose a novel GRACE with a Laplacian Smoothing learning framework that exploits contextual features in both temporal and spatial domains for robust DeepFake video detection under noisy face sequences. To the best of our knowledge, this is the first work to address the issue of unreliable face sequences for DeepFake video detection.
@@
  • We introduce a Feature Entanglement (FE) mechanism to construct an affinity matrix that mixes the spatiotemporal features together, ensuring each node contains at least one feature from valid face images. This approach effectively mitigates the impact of invalid facial images in the noisy face sequence.
@@
  • We propose a GL smoothing regularizer in the GCN to filter the noisy nodes further and improve the performance of DeepFake video detection. Comprehensive experiments demonstrate that our method achieves state-of-the-art performance, especially under challenging scenarios with unreliable and noisy face sequences.
@@
The rest of this paper is organized as follows. Section 2 presents the proposed GRACE architecture design. In Section 3, the superiority of GRACE over benchmark methods is experimentally demonstrated. Finally, conclusions are drawn in Section 4.
@@ 2.1 Overview of the Proposed Method
The proposed GRACE’s flowchart is illustrated in Fig. 2. First, a face detector extracts facial images from each video frame. A CNN-based backbone network extracts high-level semantic features from the spatial domain of the acquired facial parts, as displayed in the center of 2. Using the extracted spatial features, the spatial and temporal representations at frame $n$ and location $(i,j)$ across all feature maps ${\bm{X}}\in\mathbb{R}^{d\times c}$ can be obtained for the face sequence, where $d=N\times w\times h$, potentially including partially invalid faces. This feature representation captures feature responses at frame $n$ and location $(i,j)$ across all frames, thereby integrating temporal information, as shown in Fig. 2.
@@
To augment the correlation between the spatial and temporal feature representation ${\bm{X}}$ acquired in the previous step, we introduce a novel Feature Entanglement (FE) with sparse constraint, denoted as ${\bm{X}}_{\textrm{FE}}=G_{\textrm{FE}}({\bm{X}})\in\mathbb{R}^{d\times d}$, which carefully embeds both temporal and spatial features into its graph representation by affinity matrix ${\bm{X}}_{\textrm{FE}}$ from original feature ${\bm{X}}$. In highly noisy face sequences, the number of invalid faces could be more than that of valid ones. Therefore, the essential features could be relatively fewer, motivating us to introduce the sparsity constraint into our GRACE to focus on those essential features. Then, to efficiently discern the importance of the graph representation ${\bm{X}}_{\textrm{FE}}$, we introduce the GCN to capture the contextual features between nodes (spatiotemporal features) in ${\bm{X}}$. To further remove the noisy nodes from the original ${\bm{X}}_{\text{FE}}$, Graph Laplacian is judiciously adapted to each layer of the GCN for better performance under noisy face sequences. Finally, a softmax classifier is connected to the outcome of GCN to evaluate the authenticity of the supplied facial parts.
2.2 Feature Entanglement with Sparse Constraint
@@
We develop a method inspired by spatiotemporal feature extraction [44][45]. Traditional CNNs serve as the backbone network to obtain the spatial feature representation ${\bm{X}}^{n}\in\mathbb{R}^{c\times h\times w}$ for each frame ${\bm{y}}_{d}^{n}$ of the video. Assuming that the size of the extracted feature map is $c\times h\times w$, the spatial feature representation of a specific video at location $(i,j)$ via the backbone network can be vectorized into ${\bm{x}}_{n,i,j}=[x_{(}n,i,j)^{1},x_{(}n,i,j)^{2},...,x_{(}n,i,j)^{c}]\in\mathbb{R}^{c\times 1}$, where $c$ is the number of channels in the extracted spatial feature map and $i=1,...,w$, $j=1,...,h$. Then, let $d=Nwh$, we create the feature context ${\bm{X}}\in\mathbb{R}^{d\times c}$ based on location-wise feature concatenation, as follows:
${\bm{X}}=[$ $({\bm{x}}^{1}_{1,1,1},\ {\bm{x}}^{2}_{1,1,1},\ ...,\ {\bm{x}}^{c}_{1,1,1});$ (1)
@@
where ${\bm{X}}$ represents the spatiotemporal feature.
@@
The efficient extraction of joint spatial and temporal feature representations from ${\bm{X}}$ necessitates addressing potential inefficiencies linked to considerable distances between the $n$-th and $m$-th feature vectors, ${\bm{x}}^{n}$ and ${\bm{x}}^{m}$, especially when $m$ and $n$ are significantly apart. Directing adopting the Transformer [44][46] could lead to large computational complexity since the number of tokens could be large enough to have a long-range correlation.
@@
To address these issues and efficiently exploit joint spatial and temporal feature representations, we propose a novel concept, i.e., FE, which is exactly an affinity matrix, creating a graph representation that encapsulates the relationships between nodes—each node representing entangled spatial and temporal features based on edges. Every element (node) in ${\bm{X}}_{\textrm{FE}}$ intertwines spatial and temporal information, thereby facilitating graph neural networks to learn the relationships among spatiotemporal features more effectively, in which the features of the first and last frames might still have a link. The feature entanglement ${\bm{X}}_{\textrm{FE}}$ is defined as follows:
$\mathbf{X}_{\text{FE}}=G_{\textrm{FE}}({\bm{X}})=\mathbf{X}\mathbf{X}^{T},$ (2)
where $\mathbf{X}_{\text{FE}}\in\mathbb{R}^{d\times d}$ is the affinity matrix obtained through feature entanglement. Consider, for instance, the feature element $\mathbf{X}_{\text{FE}}({1,c})$, which is computed by taking the inner product of the first and the $c$-th row of $\mathbf{X}$. This element encapsulates the correlation between the spatial features at location $(1,1)$ across all frames and the spatial features at location $(w,h)$ in the first frame. By constructing the affinity matrix $\mathbf{X}_{\text{FE}}$ in this manner, we effectively capture the spatiotemporal dependencies within the feature representation. However, learning discriminative features from $\mathbf{X}_{\text{FE}}$ can be challenging, particularly in the presence of highly noisy face sequences, where a significant portion of the facial images may be invalid. To address this issue, we introduce a sparsity constraint as a regularization term in the learning objective, formulated as the $\ell_{1}$ norm of $\mathbf{X}$, denoted by $|\mathbf{X}|_{1}$. By enforcing sparsity on $\mathbf{X}$, we encourage the model to focus on the most informative features, thereby enhancing its robustness to noisy face sequences and improving its effectiveness in DeepFake video detection.
2.3 Proposed GCN with Graph Laplacian
In the previous subsection, we introduced the overall pipeline of our proposed GRACE method. A crucial component of GRACE is the FE with sparse constraint, which aims to efficiently exploit joint spatial and temporal feature representations. In this subsection, we delve into the details of FE and discuss how it addresses the challenges associated with unreliable face sequences in DeepFake detection.

### v3

@@ Abstract
As DeepFake video manipulation techniques escalate, posing profound threats, the urgent need to develop efficient detection strategies is underscored. However, one particular issue lies with facial images being mis-detected, often originating from degraded videos or adversarial attacks, leading to unexpected temporal artifacts on spatial domain that can undermine the efficacy of DeepFake video detection techniques. This paper introduces a novel method for robust DeepFake video detection, harnessing the power of the proposed Graph-Regularized Attentive Convolutional Entanglement (GRACE) based on the graph convolutional network to address the aforementioned challenges. First, conventional convolution neural networks are deployed to perform spatiotemporal features for the entire video. Then, the spatial and temporal features are mutually entangled by constructing a graph with sparsity constraint, enforcing essential features of valid face images in the noisy face sequences remaining, thus augmenting stability and performance for DeepFake video detection. Furthermore, the Graph Laplacian smoothing prior regularization is proposed in the graph convolutional network to suppress the noise pattern in the feature space to further improve the performance and robustness. Comprehensive experiments are conducted to illustrate that our proposed method delivers state-of-the-art performance in DeepFake video detection under noisy and unreliable face sequences. The source code is available at https://github.com/ming053l/GRACE.
@@ 1 Introduction
DeepFake image and video manipulation techniques have emerged as the most well-known forgery generation applications, with far-reaching impacts on numerous individuals. Generally, facial manipulation schemes can be classified into four categories: 1) entire face synthesis, 2) attribute manipulation, 3) identity swap, and 4) expression swap. Identity swap schemes have the most significant impact as they can be used to fabricate fake news targeting specific politicians. Many DeepFake detection techniques focus on identifying such fake videos using supervised learning methods with a pre-collected large-scale training set [5, 6, 7, 8, 9, 10, 11, 12, 13, 14].
@@
Figure 1: Example of the detected faces from two videos using RetinaFace (top) and Dlib (bottom). The sampled consecutive frames may not accurately detect faces, such noisy sequence may result in inconsistent spatial semantic features extracted by the feature extractor, which in turn makes practical application challenging.
@@
Even video compression could reduce the detection rate of the face detectors. We randomly select 200 videos from the FF++ dataset [1], featuring varying compression ratios (raw, c23, c40), and extract 16 frames from each video for face detection analysis. We employ state-of-the-art face detection tools, such as RetinaFace [25] and Dlib [26], to substantiate our observations, as shown in the Figure 1. Table I presents the face detection outcomes for the sampled videos. Notably, the predicted probability of 8 and 2 videos using RetinaFace [25] and Dlib [26] falls within the $[0,0.33]$ range, respectively, implying there are 8 and 2 facial images are mis-detected. Additionally, for uncompressed videos, 19 videos exhibit accuracy lower than $66\%$, highlighting the imperfections of face detectors. The question is raised: Could the current DeepFake detection methods be robust to such noisy face sequences? The answer is negative. We simply replace the $40\%$ facial images with background ones for the testing set of FF++ [1] with raw setting and evaluate the performance using Xception [6]. Unsurprisingly, the accuracy dropped significantly after replacement. An effective solution to deal with the issues raised by noisy face sequences for DeepFake video detection is highly desired.
Figure 2: Flowchart of the proposed GRACE for robust DeepFake video detection. In practice, face detection results are not always effective. Invalid faces may significantly outnumber valid faces, limiting the DeepFake detection methods. We leverage the proposed Feature Entanglement technique to embed the spatio-temporal features. Graph Laplacian Smoothing Prior Regularization and Sparsity Constraint to smooth and suppress these noisy nodes.
@@
In light of the escalating threat posed by various malicious attacks on face detectors that aim to undermine their reliability, this paper presents a pioneering Graph-Regularized Attentive Convolutional Entanglement (GRACE) with Laplacian Smoothing learning approach. GRACE leverages contextual features in both spatial and temporal domains to effectively detect DeepFake videos under noisy face sequences. By employing the proposed Feature Entanglement (FE) technique, an affinity matrix is constructed to amalgamate the spatiotemporal features, ensuring that each node possesses at least one feature descriptor originating from valid face images. Ultimately, Graph Laplacian Smoothing Prior regularization (GLSPR) is ingeniously integrated into the Graph Convolutional Network (GCN) to further suppress noisy nodes. Meanwhile, we meticulously incorporate Sparsity Constraint (SC) into network optimization term to prioritize the features of valid face images within the noisy face sequence, thereby significantly enhancing the performance of DeepFake video detection. The main contributions of this paper are three-fold:
@@
  • We propose a novel GRACE, which exploits contextual features in both spatial and temporal domains for robust DeepFake video detection under noisy face sequences. To the best of our knowledge, this is the first work to address the issue of unreliable face sequences for DeepFake video detection.
@@
  • We introduce a FE mechanism to construct an affinity matrix that mixes the spatiotemporal features together, ensuring the efficacious representation learning of GCN. This approach effectively mitigates the impact of invalid facial images in the noisy face sequence.
@@
  • We propose a GLSPR and SC to filter the noisy nodes further and improve the performance of DeepFake video detection. Comprehensive experiments demonstrate that our method achieves state-of-the-art performance, especially under challenging scenarios with unreliable and noisy face sequences.
@@
The rest of this paper is organized as follows. Section 2 introduces the related works in DeepFake detection. Section 3 presents the proposed GRACE architecture design. In Section 4, the superiority of GRACE over benchmark methods is experimentally demonstrated. Finally, conclusions are drawn in Section 5.
2 Related Works
2.1 General DeepFake Video Detection
In the domain of DeepFake video detection, numerous sophisticated approaches have recently emerged [27][28][24][22][21][29] [30][31][32][33]. On the one hand, these methods extend DeepFake image detection techniques by averaging the predictions of individual frames to assess a video’s authenticity [24][31][32][33]. On the other hand, the temporal inconsistency feature is exploited for DeepFake video classification using supervised learning approaches, as demonstrated in [21][29][27][28]. Recently, several advanced techniques have been proposed to enhance DeepFake video detection performance. To address the generalizability issue, semi-supervised learning is considered in [34, 35, 36] to capture common fake features from selected representative GANs, assuming that most GANs might share similar identifiable clues. Pairwise learning is then employed to learn these common features from the training set, improving generalizability for DeepFake image detection [35, 34]. CORE [37] introduces a novel approach for learning consistent representations across different frames, while RECCE [38] employs a reconstruction-classification learning scheme to capture more discriminative features. DFIL [39] proposes an incremental learning framework that exploits domain-invariant forgery clues to improve generalization ability. TALL-Swin [40] utilizes a thumbnail layout and Swin Transformer to learn robust spatiotemporal features for DeepFake detection. UCF [41] focuses on uncovering common features shared by different manipulation techniques to enhance generalizability. DFGaze [42] uses the gaze analysis to analyze the gaze features of face video frames, and then applies a spatial-temporal feature aggregator to realize authenticity classification.
2.2 External Reference DeepFake Video Detection
As Face X-ray [24] demonstrates superior performance and robustness, recent DeepFake video detection techniques concentrate on uncovering more reliable signatures produced by GANs to enhance detection performance. Specifically, state-of-the-art DeepFake video detection primarily focuses on exploiting various priors, such as bio-informatics clues [32], facial warping artifacts [33], and noise patterns [24]. FakeCatcher [32] utilizes a novel bio-feature, Photoplethysmography (PPG) response, to differentiate DeepFake videos, as real and fake videos exhibit distinct PPG features. A critical limitation is the necessity for high-resolution videos and images to effectively capture PPG cues. Moreover, The study in [33] investigates warping artifacts at the boundaries of DeepFake videos, which arise due to the limited resolution of synthesized facial components. Concurrently, The study in [27] introduces a method utilized deep Laplacian of Gaussian (LoG) and the loss of isolated manipulated faces to bolster the generalizability of DeepFake video tasks.
2.3 Robust DeepFake Video Detection
Recent research has concentrated on developing robust DeepFake detection models for compressed videos, as seen in [21][23][43]. The frequency component analysis method is employed to uncover intrinsic features and enhance performance under compression settings [23]. However, frequency-aware features may prove ineffective under high compression (with high-frequency reduction) or noisy conditions (with high-frequency amplification). The $F^{3}$-net [43] selects two complementary frequency bands as clues, devising a novel network to learn frequency-aware features that reveal subtle forgery artifacts. Recent studies [44][45][46] have highlighted the vulnerability of DeepFake detectors to adversarial perturbations. Therefore, adversarial defense with DeepFake detection, such as [47], have been attracting recently. The study in [47] demonstrates an effective solution by leveraging statistical inference on CNNs to achieve better robustness to adversarial examples.
@@ 3.1 Overview of the Proposed Method
The proposed GRACE’s flowchart is illustrated in Figure 2, which is the video-based approach. Previous research based on frequency analysis methods [43] has suggested that high-frequency information could be critical in DeepFake detection. However, in practical scenarios, the detection process can be complicated by adversarial attacks or rapid object movements, leading to erroneous face detections and resulting in network vulnerability.
@@
First, a face detector extracts facial images from each video frame. A CNN-based backbone network extracts high-level semantic features from the spatial domain of the acquired facial parts. Using the extracted spatial features, the spatial and temporal representations at frame $n$ and location $(i,j)$ across all feature maps $\mathbf{X}\in\mathbb{R}^{d\times c}$ can be obtained for the face sequence, where $d=N\times h\times w$, potentially including partially invalid faces. This feature representation captures feature responses at frame $n$ and location $(i,j)$ across all frames, thereby integrating temporal information, as shown in Figure 2.
@@
To augment the correlation between the spatial and temporal feature representation $\mathbf{X}$ acquired in the previous step, we introduce a novel Feature Entanglement, denoted as $\mathbf{X}_{\textrm{FE}}=G_{\textrm{FE}}(\mathbf{X})\in\mathbb{R}^{d\times d}$, which carefully embeds both temporal and spatial features into its graph representation by affinity matrix $\mathbf{X}_{\textrm{FE}}$ from original feature $\mathbf{X}$. In highly noisy face sequences, the number of invalid faces could be more than that of valid ones. Therefore, the essential features could be relatively fewer. To efficiently discern the importance of the graph representation $\mathbf{X}_{\textrm{FE}}$, we introduce the GCN to capture the contextual features between nodes (spatiotemporal features) in $\mathbf{X}$. To further remove the noisy nodes from the original $\mathbf{X}_{\text{FE}}$, GLSPR is judiciously adapted to each layer of the GCN for better performance under noisy face sequences. Besides, we introduce SC into the objective function to allow our GRACE method to focus on these essential features. Finally, a softmax classifier is connected to the outcome of the GCN to evaluate the authenticity of the supplied facial parts.
3.2 Feature Entanglement
@@
Traditional CNNs serve as the backbone network to obtain the spatial feature representation for each frame of the video. Assuming that the size of the extracted feature map is $c\times h\times w$, the spatial feature representation of a specific video at location $(i,j)$ via the backbone network can be vectorized into $\mathbf{x}_{n,i,j}=[x_{n,i,j}^{1},x_{n,i,j}^{2},...,x_{n,i,j}^{c}]\in\mathbb{R}^{c\times 1}$, where $c$ is the number of channels in the extracted spatial feature map, $i=1,...,h$, $j=1,...,w$, and $n=1,...,N$, with $N$ being the total number of frames. Then, let $d=N\times h\times w$, we create the feature context $\mathbf{X}\in\mathbb{R}^{d\times c}$ based on location-wise feature concatenation, as follows:
$\mathbf{X}=[$ $({\bm{x}}^{1}_{1,1,1},\ {\bm{x}}^{2}_{1,1,1},\ ...,\ {\bm{x}}^{c}_{1,1,1});$ (1)
@@
where $\mathbf{X}$ represents the spatiotemporal feature.
@@
The joint spatial and temporal feature representations from $\mathbf{X}$ necessitates addressing potential inefficiencies linked to considerable distances between the $n$-th and $m$-th feature vectors, ${\bm{x}}^{n}$ and ${\bm{x}}^{m}$, especially when $m$ and $n$ are significantly apart. Directing adopting the Transformer [48][49] could lead to large computational complexity since the number of tokens could be large enough to have a long-range correlation.
@@
To address these issues and efficiently exploit joint spatial and temporal feature representations, we propose a novel concept, i.e., Feature Entanglement, which is exactly an affinity matrix, creating a graph representation that encapsulates the relationships between nodes—each node representing entangled spatial and temporal features based on edges. Every element (node) in $\mathbf{X_{\textrm{FE}}}$ intertwines spatial and temporal information, thereby facilitating GCN to learn the relationships among spatiotemporal features more effectively, in which the features of the first and last frames might still have a link. The feature entanglement $\mathbf{X}_{\textrm{FE}}$ is defined as follows:
$\mathbf{X}_{\text{FE}}=G_{\textrm{FE}}(\mathbf{X})=\mathbf{X}\mathbf{X}^{T},$ (2)
where $\mathbf{X}_{\text{FE}}\in\mathbb{R}^{d\times d}$ is the affinity matrix obtained through feature entanglement. Consider, for instance, the feature element $\mathbf{X}_{\text{FE}}({1,c})$, which is computed by taking the inner product of the first and the $c$-th row of $\mathbf{X}$. This element encapsulates the correlation between the spatial features at location $(1,1)$ across all frames and the spatial features at location $(h,w)$ in the first frame, as shown in Figure 2.
3.3 GCN with Graph Laplacian Smoothing